OpenStudy (anonymous):

can someone help me write the equation of the line which passes through (2, 1) and is perpendicular to x = –2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

slopes of perpendicular lines are negative reciprocals of each other

OpenStudy (anonymous):

so the slope is x=2

OpenStudy (anonymous):

Am not sure about how to write the equation

OpenStudy (anonymous):

it gave me y=1x-3 so tha the equation they are asking me for?

OpenStudy (anonymous):

check ur calculation

OpenStudy (anonymous):

I really see anything wrong??

OpenStudy (anonymous):

dont see anything wrong :/

OpenStudy (anonymous):

the graph of x=-2 will be a vertical line the line perpendicular to it will be a horizontal line horizontal lines have slope = 0 so the eqn will be y-1=0(x-2) so y = 1

OpenStudy (anonymous):

oh the 0 is what i had wrong , thank you

OpenStudy (anonymous):

!!